Circular No. 7129

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

S/1998 (45) 1 W. J. Merline, Southwest Research Institute, Boulder, reports for a

large collaboration (including L. M. Close, C. Dumas, C. R. Chapman, F. Roddier, F. Menard, D. C. Slater, G. Duvert, C. Shelton, and T. Morgan) the discovery of a satellite of (45) Eugenia on 1998 Nov. 1.5 UT from H-, J-, and K′-band direct imaging with the 3.6-m Canada-France-Hawaii Telescope (+ PUEO adaptive-optics system) on Mauna Kea. The satellite is ∼ 6 mag fainter than (45) and was tracked intermittently on five nights over a 10-day span. Merline et al. derive a near-circular orbit with a period of ∼ 4.7 days, inclined to the line-of-sight by ∼ 45^{o}, with a maximum elongation of ∼ 0′′.8. The satellite was recovered with the CFHT on 1999 Jan. 4. An occultation around Mar. 27.13 of the star CMC 804951 (mag 11.9) by (45) was predicted by D. W. Dunham, but this is likely to be visible only near the earth’s north-polar regions (see updates

at http://members.home.net/dega/astchart.htm).

U SCORPII P. Bonifacio, P. Molaro, and P. Selvelli, Astronomical Observatory,

Trieste, write: “Spectra (range 400–665 nm, resolution 0.34 nm) of U Sco, obtained on Mar. 16.4 and 17.4 UT with the ESO 3.5-m New Technology Telescope (+ EMMI spectrograph), show composite structure in the hydrogen emission lines. The central component falls at the rest wavelength, with the two satellite emissions shifted by velocities of +1580 and −1760 km/s (as measured in Hα); all three components are well resolved (FWHM 600 km/s). The two emissions on either side of Hα are not attributable to [N II] (654.8 and 658.4 nm), since the triplet-like structure is seen in the other Balmer lines observable in our spectra (up to Hδ). A similar structure is present also in the He II (454.2, 468.6, 541.3 nm) emission lines. Two otherwise unidentified emission lines at 584.4 and 590.8 nm are likely the symmetrically displaced components of the He I 587.6-nm line, whose central component is absent. Such peculiar structure in both the H and He recombination lines was not reported at the 1979 and 1987 outbursts. The Hα FWZI is 5600 km/s, showing a decrement from the value reported on IAUC 7118. The relative emission intensities of the line components changed between the two nights, with a strengthening of the central one. The 464.0-nm complex is present as a wide and unresolved blend. Narrow emissions of O I (the nebular 630.0- and 636.3-nm and the auroral O I 557.7-nm transitions) are present, marking the onset of the nebular phase.”

1999 March 20  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green